# Mitigating Negative Side Effects in Multi-Agent Systems Using Blame Assignment

### v1

@@ 8 Summary and Future Work
This paper formalizes the problem of mitigating NSEs in cooperative multi-agent settings as a decentralized, bi-objective problem. The agents’ assigned tasks follow transition and reward independence. The agents produce no NSE when operating in isolation but their joint actions produce NSE and incur a joint penalty. The agents have no prior knowledge about the side effects of their joint actions. We present a metareasoning approach to detect NSEs and update agent policies in a decentralized manner, by decomposing the joint NSE penalty into individual penalties. Our algorithm, RECON, uses a counterfactual-based blame attribution to estimate each agent’s contribution towards the joint penalty. Our experiments demonstrate the effectiveness of our approach in mitigating NSEs, by updating the policies of a subset of agents.
Our framework currently supports Dec-MDPs with transition and reward independence. In the future, we aim to relax this assumption and extend our approach to settings with tightly coupled task assignment. Another interesting future direction is to exploit the agent dependencies and interactions for task completion to mitigate NSEs, by leveraging their complementary skills.

### v2

@@ VI SUMMARY AND FUTURE WORK
This paper formalizes the problem of mitigating NSEs in cooperative multi-agent settings as a decentralized, bi-objective problem. The agents’ assigned tasks follow transition and reward independence. The agents produce no NSE when operating in isolation but their joint actions produce NSE and incur a joint penalty. We present a metareasoning approach to detect NSEs and update agent policies in a decentralized manner, by decomposing the joint NSE penalty into individual penalties. Our algorithm, RECON, uses a counterfactual-based blame attribution to estimate each agent’s contribution towards the joint penalty. Our experiments demonstrate the effectiveness of our approach in mitigating NSEs. Our framework currently supports Dec-MDPs with transition and reward independence. In the future, we aim to relax this assumption and extend our approach to settings with tightly coupled task assignment.